OpenStudy (gabylovesyou):

Last year, Scott had $20,000 to invest. He invested some of it in an account that paid 6% simple interest per year, and he invested the rest in an account that paid 10% simple interest per year. After one year, he received a total of $1760 in interest. How much did he invest in each account?

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get with this?

jimthompson5910 (jim_thompson5910):

were you able to get started at all?

OpenStudy (gabylovesyou):

no .-.

jimthompson5910 (jim_thompson5910):

ok let's say we have 2 unknown variables x and y x = amount of money invested at 6% y = amount of money invested at 10% let me know when you're ready for the next step

OpenStudy (gabylovesyou):

ok ready

jimthompson5910 (jim_thompson5910):

we can only invest money at 6% or 10%. There are no other choices. If we have $20,000 to invest, then that means x and y must add to 20,000 ie x+y = 20,000 do you agree?

OpenStudy (gabylovesyou):

yes

jimthompson5910 (jim_thompson5910):

now let's isolate a variable. Let's say we isolate y how would we get y all by itself? what must we do to both sides?

OpenStudy (gabylovesyou):

subtract x ?

jimthompson5910 (jim_thompson5910):

yep subtract x from both sides

jimthompson5910 (jim_thompson5910):

x+y = 20,000 x+y-x = 20,000-x y = 20,000 - x

OpenStudy (gabylovesyou):

yes

jimthompson5910 (jim_thompson5910):

this will be used later. So we'll come back to it

jimthompson5910 (jim_thompson5910):

Scott invested x dollars at 6% How much money does he earn just from this account? Any idea?

OpenStudy (gabylovesyou):

does it have to equAL 20000 ?

jimthompson5910 (jim_thompson5910):

ignore that part for now

jimthompson5910 (jim_thompson5910):

if you had $100 for example, and you earn 6% interest, how much money have you earned?

OpenStudy (gabylovesyou):

ummm.

jimthompson5910 (jim_thompson5910):

interest = principal*rate*time principal = amount you invest rate = interest rate in decimal form time = time in years (in this case, we'll make the time = 1 to be 1 year)

OpenStudy (gabylovesyou):

100 * 6%*1

jimthompson5910 (jim_thompson5910):

6% = 6/100 = 0.06

jimthompson5910 (jim_thompson5910):

so you'd have 100*0.06*1 = ????

OpenStudy (gabylovesyou):

6

jimthompson5910 (jim_thompson5910):

yes, so now instead of $100 we don't know what we invest at 6% so we'd have x*0.06*1 = ????

OpenStudy (gabylovesyou):

.06x

jimthompson5910 (jim_thompson5910):

that means Scott earned 0.06x dollars so far (from the 6% investment) x is some unknown number

OpenStudy (gabylovesyou):

ok

jimthompson5910 (jim_thompson5910):

yes, now Scott also invested y dollars (y is some unknown number) at 10% how much has Scott earned from just the 10% investment?

OpenStudy (gabylovesyou):

x*.10*1

jimthompson5910 (jim_thompson5910):

which simplifies to what?

jimthompson5910 (jim_thompson5910):

btw use y instead of x

OpenStudy (gabylovesyou):

.10y

jimthompson5910 (jim_thompson5910):

0.10y yes

jimthompson5910 (jim_thompson5910):

Scott earned 0.06x dollars from the 6% account he also earned 0.10y dollars from the 10% account in total, he earned 0.06x+0.10y dollars do you agree?

OpenStudy (gabylovesyou):

yes

jimthompson5910 (jim_thompson5910):

we're also told that `After one year, he received a total of $1760` so that must mean the expression I just got is equal to 1760, so, 0.06x+0.10y = 1760 make sense?

OpenStudy (gabylovesyou):

yes

jimthompson5910 (jim_thompson5910):

we now go back to `y = 20,000 - x` replace every copy of 'y' in the the second equation with 20,000-x so we'd go from 0.06x+0.10y = 1760 to 0.06x+0.10(20,000 - x) = 1760 Don't forget about the parenthesis. From here, you would solve for x

OpenStudy (gabylovesyou):

19999.719

OpenStudy (gabylovesyou):

i need the amount separately. hm he invested in one account and then the other,

jimthompson5910 (jim_thompson5910):

how did you get 19999.719 ?

OpenStudy (gabylovesyou):

mathway cx

jimthompson5910 (jim_thompson5910):

it's better to do it by hand on pencil and paper

jimthompson5910 (jim_thompson5910):

what would be the first step to solve `0.06x+0.10(20,000 - x) = 1760` ?

OpenStudy (gabylovesyou):

ok idk how to do this.. i have so many more problemsss .-.

jimthompson5910 (jim_thompson5910):

distribute the 0.10 through. What do you get?